OpenStudy (anonymous):

Is this a trick question? -(1/3)^0 and (-8)^0 I just got asked this in a review work sheet and I'm wondering if there is something I may have skipped in my lessons? To the power of 0 doesn't just mean that the number stays the same? Help please! Thank you

OpenStudy (nikato):

no. any number to the power of 0 is 1

OpenStudy (tkhunny):

10^3 = 1000 10^2 = 100 10^1 = 10 10^0 = 10? No.

OpenStudy (anonymous):

I feel stupid, yeah that's what I meant...not the same number. Like 7 times 0 = 0 and such

OpenStudy (anonymous):

but if that's the case then it is sort of a trick question then? The answer to both is just 0?

OpenStudy (tkhunny):

10^0 = 1 -- That is not zero.

OpenStudy (anonymous):

So it is different when it's to the power of rather than multiplication. That's what I wanted to know. Thank you.

OpenStudy (tkhunny):

We do try to make things consistent. Try a few values around where you want to explore and imagine what "consistent" behavior might be. That's all I did. 10^3 = 1000 10^2 = 100 10^1 = 10 10^0 = 1 10^(-1) = 1/10 10*3 = 30 10*2 = 20 10*1 = 10 10*0 = 0 10*(-1) = -10 They should all make sense at the same time.

OpenStudy (anonymous):

Thank you so much for your thorough explanation :)